Use the data below obtained when 5.00 g of zinc were reacted with various amounts of 2.0 M HCl and the graph to answer the questions that follow. Zn (s) 2 HCl (aq) ZnCl2 (aq) H2 (g) Mass Zn (g) Volume 2.0 M Mass H H2 Produced when 5.00g zn Reacted With HCI (mL) formed (g) HCI 5.00 0.0 5.00 20.0 0040 0.200 150 5.00 40.0 5.00 60.0 5.00 80.0 0.154 0.100 0.050 0.000 100.0 20.0 40.0 600 80.0 20.0 140.0 Volume 2.00M HCI Gmi) 5.00 120.0 0.154 1. Which reactant is limiting the amount of H2 produced in the region of the graph labeled A? Region B? 2. How many grams of hydrogen gas can be produced if 200.0 mL of 2.0 M HCl are reacted with 5.00 grams of zinc? 3. How many grams of Zn are required to completely react with 200.0 mL of 2.0 M HCl? How many grams of H2 can be produced? (use back of this page if necessary)
Answer #1

For given information,

Zn (s) 2 HCI (aq)-> ZnCl2 (aq) H2 (g) Mass Zn (g)Volume 2.0 MMass H2 HCl (ml 0.0 20.0 40.0 60.0 80.0 100.0 120.0 H2 Produced

The reaction is

Zn(s) + 2HCl(aq) = ZnCl2(aq) + H2(g)

According to stoichiometry, 1 mole of Zn reacts with 2 moles of HCl to form each moles of ZnCl2 and H2.

Molecular mass of Zn = 65.4 g/mol

Zinc reacting = 5 gm

moles of Zn reacting = Wt / mol. Wt = 5/65.4 = 0.076 mols

Molarity of HCl is 2.0M i.e. 2 mols/lit

Moles of HCl = Molarity * Volume in litre

Volume of 2M HCl (ml) Moles of HCl
0 0
20 0.04
40 0.08
60 0.12
80 0.16
100 0.2
120 0.24

For 5g of Zn,i.e. 0.076 mols, according to stoichiometry, 0.152 mols of HCl are required to react.

Hence, untill the moles of HCl are less than 0.152, for the reaction, HCl is the limiting reagent. i.e. upto 80ml of HCl

and After the moles of HCl are higher than 0.152, for the reaction, Zn is the limiting reagent.i.e, from 80ml of HCl onwards.

In graph, the region A has HCl as limiting reactant and the region B has Zn as limiting reactant.

2)

If 200ml of 2.0M HCl is reacted with 5g of Zn,

200ml of 2.0M HCl has 0.4 mols of HCl

and 5g of Zn has 0.076 moles of Zn.

According to stoichiometry, here, Zn is the limiting reactant.

hence 0.076 moles of Zn would produce 0.076 moles of H2.

Molecular weight of H2 = 2 gm/mol

mass of H2 produced = mol. wt * no. of moles

mass of H2 produced = 0.152g

3)

200ml of 2.0M of HCl has 0.4 moles of HCl.

According to stoichiometry, 0.4 moles of HCl require 0.2 moles of Zn to react completely.

mass of 0.2 moles of Zn = Mol. Wt * no. of moles

Mass of 0.2 moles of Zn = 65.4*0.2

Mass of 0.2 moles of Zn = 13.08g

0.4 moles of HCl reacts completely with 13.08g of Zn to produce 0.2 moles of H2

Mass of 0.2 moles of H2 = Mol.wt * no. of moles

Mass of 0.2 moles of H2 = 2*0.2

Mass of 0.2 moles of H2 = 0.4g

Hence 200ml of 2.0M HCl would require 13.08g of Zn to react completely to produce 0.4g of H2 gas